# Vevo2: Bridging Controllable Speech and Singing Voice Generation via Unified Prosody Learning

### v1

@@ 3.1 Prosody Tokenizer
Architecture The prosody tokenizer consists of five main components: Downsampling Layer, Encoder, Vector Quantization (VQ), Decoder, and Upsampling Layer. Formally, given the codebook $\bm{E}=[\bm{e}_{1},\bm{e}_{2},\dots,\bm{e}_{K}]$ with vocabulary size $K$, and taking chromagram features $\bm{x}$ as input, the reconstructed output $\hat{\bm{x}}$ is obtained after passing through these five modules:
@@
We follow Yu et al. (2022); Kumar et al. (2023); Wang et al. (2025a) and adopt factorized (i.e., a low-dimensional code-lookup space) and L2-normalized strategies for VQ learning. We extract 50 Hz chromagram features from raw audio and apply an 8x downsampling, resulting in a prosody tokenizer operating at 6.25 Hz (56.25 bps).
3.2 Content-Style Tokenizer
@@
Motivation We follow the main ideas from recent works in controllable speech generation Zhang et al. (2025a); Du et al. (2024a), and aim to develop a content-style tokenizer that can encode linguistic content, prosody, and style, while also achieving timbre disentanglement. Specially, to unify both speech and singing voice, beyond encoding linguistic content and style, as in typical speech generation tasks Zhang et al. (2025a); Du et al. (2024a); Anastassiou et al. (2024a); Wang et al. (2025a), we also require the tokenizer to capture melody and fine-grained style information (e.g., F0 and energy details) in singing voice. Motivated by this, when designing our VQ-VAE tokenizer, we not only use hidden features from automatic speech recognition (ASR) models—specifically, from Whisper Radford et al. (2023)—as reconstruction objectives, but also incorporate chromagram features as an additional reconstruction objective. This compensates for the limitations of Whisper features in modeling melody in singing voice Zhang et al. (2024e). Furthermore, to facilitate the downstream content-style modeling stage, we design the content-style tokenizer to operate at a low frame rate (specifically, 12.5 Hz in our study, which is lower than the commonly used 25 or 50 Hz Anastassiou et al. (2024a); Zhang et al. (2025a); Wang et al. (2025a); Du et al. (2024a, b)). This reduces the sequence length for the AR model and thus alleviates its learning burden.
@@
Architecture The architecture of the content-style tokenizer is similar to the proposed prosody tokenizer. The only difference lies in the input and reconstruction objectives incorporate both chromagram features and Whisper features. We extract 50 Hz chromagram features and 50 Hz Whisper features (which is the encoder output of the pre-trained Whisper model Radford et al. (2023)) and apply a 4x downsampling, resulting in a content-style tokenizer operating at 12.5 Hz (175 bps).
3.3 Pre-training
Content-Style Modeling For content-style modeling, the common approach in speech generation works is to learn a mapping from text tokens to content-style tokens (typically using an auto-regressive Borsos et al. (2023); Anastassiou et al. (2024a); Du et al. (2024a); Zhang et al. (2025a) or masked generative model Wang et al. (2025a)). We refer to this as Implicit Prosody Learning (IPL), since the model implicitly supplements prosody information for the mapping through in-context learning. However, for singing voice generation, it is often desirable to explicitly control the prosody (i.e., melody) of the generated output Umbert et al. (2015); Liu et al. (2022); Zhang et al. (2024c); Zhao et al. (2025). To address it, we propose an Explicit Prosody Learning (EPL) strategy, where prosody tokens are provided as explicit input, allowing the model to predict content-style tokens from both text and prosody tokens. Notably, during training, we do not restrict IPL to speech data and EPL to singing voice data. Instead, we treat all samples equally: for each training sample, we randomly choose between IPL and EPL with equal probability. We assume that this design better unifies the training patterns for speech and singing voice.
During pre-training, we initialize the AR transformer with a pre-trained textual large language model (LLM), specifically Qwen2.5-0.5B Yang et al. (2024b). Given a speech or singing voice sample u{u}, we denote the prosody and content-style tokens as $\bm{Q}_{p}({u})$ and $\bm{Q}_{cs}({u})$, respectively. The input sequences to the transformer for EPL and IPL are shown in Equation 3 and Equation 4, where $\bm{I}_{ipl}$ and $\bm{I}_{epl}$ represent the text instruction prefixes for IPL and EPL (see our Appendix for more details), $\bm{T}$ denotes the text of u{u}, and $\langle|\cdot|\rangle$ indicates special tokens. We perform next token prediction only on the final segment $\langle|\text{start-of-cs}|\rangle,\bm{Q}_{cs}({u}),\langle|\text{end-of-cs}|\rangle]$:
$\bm{I}_{ipl},\bm{T},\langle|\text{start-of-cs}|\rangle,\bm{Q}_{cs}({u}),\langle|\text{end-of-cs}|\rangle].$ (3)
$\bm{I}_{epl},\bm{T},\langle|\text{start-of-p}|\rangle,\bm{Q}_{p}({u}),\langle|\text{end-of-p}|\rangle,\langle|\text{start-of-cs}|\rangle,\bm{Q}_{cs}({u}),\langle|\text{end-of-cs}|\rangle].$ (4)
During inference, when the prosodic source is not used, the inference paradigm is identical to classic zero-shot TTS Wang et al. (2023); Anastassiou et al. (2024a); Zhang et al. (2025a); Du et al. (2024a). As illustrated in the TTS task in Figure 1, the process can be represented as $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$], where the style signals in $\boldsymbol{\bm{\hat{cs}}}$ are prompted by $\boldsymbol{\bm{cs}}$ and generated through the in-context learning capabilities of the pre-trained model. Particularly, when both prosodic source and style reference are utilized as inputs, the inference paradigm takes the form $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$]. In this context, for $\boldsymbol{\bm{\hat{cs}}}$, we expect $\boldsymbol{\bm{p}}$ to provide coarse-grained prosody contour signals, while $\boldsymbol{\bm{cs}}$ contributes fine-grained style details (e.g., the vibrato patterns for singing). We will discuss this further through our experiments on style-converted singing voice conversion (Section 4.3).
Acoustic Modeling During acoustic modeling, we follow the common approach adopted in speech generation studies Bai et al. (2024); Du et al. (2024a); Zhang et al. (2025a); Zeng et al. (2025), employing a flow-matching transformer to convert content-style tokens into Mel spectrograms, conditioned on a timbre reference. After obtaining the Mel spectrogram, we use a Vocos-based vocoder Siuzdak (2024) that is pre-trained on both speech and singing voice, to generate the final waveform. We provide the detailed architecture designs in Appendix A.3.
3.4 Post-training
Motivation After pre-training, we observe that the AR model exhibits good versatility across diverse tasks (Section 4.3). Nevertheless, its stability remains suboptimal, particularly in its fundamental capabilities of text- and prosody-following, which can be further elicited and enhanced. Furthermore, in downstream applications, the model may encounter out-of-pretraining data such as instrumental sounds. For instance, in the context of SVS, the model needs to accept prosody tokens extracted from MIDI-rendered sounds for melody control (Figure 1). Motivated by the above, this study further employs post-training to enhance the model’s controllability for text and prosody, as well as to improve its robustness to diverse audio data. We propose a multi-objective post-training task that jointly aligns both intelligibility and prosody similarity (Figure 2). We believe that this joint alignment helps prevent the model from over-optimizing a single objective (as evidenced in Section 4.4).
@@ 3.4 Post-training
Intelligibility Reward To align intelligibility, we adopt a recent zero-shot TTS intelligibility preference corpus, INTP Zhang et al. (2025b), which contains diverse intelligibility preference pairs constructed through comparisons between multiple models as well as manually designed intelligibility perturbations, to train an intelligibility reward model. Specifically, we employ the commonly used Bradley-Terry (BT) approach Bradley and Terry (1952); Ouyang et al. (2022); Bai et al. (2022) for reward model training. Given an intelligibility preference dataset $\mathcal{D}=\{(t,a^{w},a^{l})\}$, where $a^{w}$ and $a^{l}$ represent the positive and negative audio outputs with respect to the target text $t$, we parametrize an intelligibility reward model $r_{\phi}(t,a)$, treat the problem as binary classification, and optimize the negative log-likelihood loss:
@@
where $r_{\phi}(t,a)$ is initialized from the pre-trained model. We follow the pre-training and use both EPL and IPL for training this reward model. During inference, we can input it a $\bm{T},\bm{Q}_{cs}]$ or $\bm{T},\bm{Q}_{p},\bm{Q}_{cs}]$ sequence to obtain an intelligibility reward. More details are provided in Appendix A.4.1.
Prosody Similarity Reward To achieve prosodic alignment, we establish the following protocol to evaluate the policy model’s prosody-following capability: (1) Suppose we have a singing voice dataset with MIDI annotations. Given a singing voice sample u{u}, we use its MIDI-rendered instrumental music to extract prosody tokens as control conditions, denoted as $\bm{Q}_{p}$; (2) Let the policy model perform a completion conditioned on $\bm{Q}_{p}$ and $\bm{T}$ (where $\bm{T}$ can be any text), and denote the generated content-style tokens as $\hat{\bm{Q}}_{cs}$; (3) Using the decoder of the content-style tokenizer (Section 3.2), we reconstruct the chromagram features $\hat{\bm{x}}$ from $\hat{\bm{Q}}_{cs}$, and compute the cosine similarity between $\hat{\bm{x}}$ and the ground-truth chromagram features $\bm{x}$ of the singing voice u{u}. [^2] Note that in step (3), we could also use the flow-matching model and vocoder to perform inference and extract $\hat{\bm{x}}$ from the predicted waveform, but for training efficiency, we adopt the current approach. In practice, our prosody prompt data includes not only the aforementioned instrumental music data, but also some of the speech and singing voice data used during pre-training. More details are provided in Appendix A.4.2.
Multi-Objective Optimization We apply the GRPO algorithm Shao et al. (2024) to optimize the proposed multi-objective alignment task. Assuming the number of completions per prompt is $K$, for each prompt, after obtaining the intelligibility rewards ($\bm{r}_{int}$) and prosody similarity rewards ($\bm{r}_{pro}$) for all $K$ completions, we apply group normalization to both rewards separately. The sum of these normalized rewards serves as the final outcome supervision:
@@
3.5 Inference-Time Control
@@
After obtaining the pre-trained (or post-trained) AR model and FM model, we can perform inference following the pipeline illustrated in Figure 1, by utilizing different combinations of text, prosodic source, style reference, and timbre reference to achieve versatile synthesis, conversion, and editing tasks for both speech and singing voice. In this section, we further explain two additional inference-time control mechanisms of Vevo2: Duration Control and Pitch Region Control.
@@
Duration Control Some Non-AR based speech generation models have often featured the ability to control the total duration of generated speech Wang et al. (2025a); Chen et al. (2025), while AR-based models are generally considered less capable of controlling output duration Borsos et al. (2023); Anastassiou et al. (2024a); Du et al. (2024a); Zhang et al. (2025a); Zhou et al. (2025). Interestingly, however, we discover that Vevo2 can effectively control the duration of generated results through the use of prosody tokens as input. This capability stems from our tokenizers’ fixed frame rates (6.25 Hz and 12.5 Hz), which maintain a consistent 2:1 ratio between content-style and prosody token sequences during training. Consequently, during inference, we can control the duration of the generated output by applying simple linear scaling transformations to the chromagram features of the prosodic source, thereby modifying the length of prosody tokens. Such control mechanism is detailed in Appendix A.5.1.
@@
Pitch Region Control For voice conversion (VC) or singing voice conversion (SVC) tasks, due to potential significant differences in timbre types Wilkins et al. (2018); Sundberg and Rossing (1990) between source and reference speakers, we often need to adjust the pitch region (i.e., apply pitch shift) of the source waveform during inference to achieve higher speaker similarity Huang et al. (2023); Zhang et al. (2024e); Sha et al. (2024). In Vevo2, we discover that pitch region control can also be achieved: by shifting the pitch of the source waveform before prosody token extraction, we can effectively control the pitch region of the generated speech or singing voice. Our experiments demonstrate that this approach is effective both when using only the FM stage and when combining AR and FM stages. Such control mechanism is detailed in Appendix A.5.2.
4 Experiments

### v2

@@ III-A Unified Speech and Singing Voice Tokenizers
The prosody tokenizer consists of five main components (Figure 2(a)): Downsampling Layer (CNN-based), Encoder, Vector Quantization (VQ), Decoder, and Upsampling Layer (CNN-based). Formally, given the codebook $\bm{E}=[\bm{e}_{1},\bm{e}_{2},\dots,\bm{e}_{K}]$ with vocabulary size $K$, and taking chromagram features $\bm{x}$ as input, the reconstructed output $\hat{\bm{x}}$ is obtained after passing through these five modules:
@@
We follow [vqgan-cv, rvqgan, maskgct] and adopt factorized (i.e., a low-dimensional code-lookup space) and L2-normalized strategies for VQ learning. We extract 50 Hz chromagram features from raw audio and apply an 8x downsampling, resulting in a prosody tokenizer operating at 6.25 Hz (56.25 bps).
@@
Content-Style Tokenizer We follow the main ideas from recent works in controllable speech generation [vevo, cosyvoice], and aim to develop a content-style tokenizer that can encode linguistic content, prosody, and style, while also achieving timbre disentanglement. Specially, to unify both speech and singing voice, beyond encoding linguistic content and style, as in typical speech generation tasks [vevo, cosyvoice, seedtts, maskgct], we also require the tokenizer to capture melody and fine-grained style information (e.g., F0 and energy details) in singing voice. Motivated by this, when designing our VQ-VAE tokenizer, we not only use hidden features from automatic speech recognition (ASR) models—specifically, from Whisper [whisper]—as reconstruction objectives, but also incorporate chromagram features as an additional reconstruction objective. This compensates for the limitations of Whisper features in modeling melody in singing voice [amphion-svc]. Furthermore, to facilitate the downstream content-style modeling stage, we design the content-style tokenizer to operate at a relatively low frame rate (specifically, 12.5 Hz in our study). This reduces the sequence length for the AR model and thus alleviates its learning burden.
@@
The architecture of the content-style tokenizer is similar to the proposed prosody tokenizer, which is displayed in Figure 2(b). The only difference lies in the input and reconstruction objectives incorporate both chromagram features and Whisper features. We extract 50 Hz chromagram features and 50 Hz Whisper features (which is the encoder output of the pre-trained Whisper model [whisper]) and apply a 4x downsampling, resulting in a content-style tokenizer operating at 12.5 Hz (175 bps).
III-B Speech-Singing Joint Training
To conduct the joint training of speech and singing voice, an essential question lies in how to unify the generation process of speech and singing within a single framework. For the acoustic modeling stage, once the proposed unified content-style tokenizer is obtained, this part becomes relatively straightforward, as both speech and singing voice can be formulated into unified content-style tokens and decoded into Mel spectrograms under the guidance of a timbre reference. However, we observe that the main challenge of speech-singing joint training arises in the content-style modeling stage, where the generation characteristics of speech and singing differ substantially. For instance, speech generation like TTS usually takes only text (and the reference) as input and does not use an explicit control over prosody [tts-book-tanxu], while singing voice generation often demands direct control over melody [singing-expression, diffsinger, tcsinger, spsinger].
@@ III-B Speech-Singing Joint Training
Motivated by this observation, we focus on the unified prosody learning and propose two different training strategies to bridge speech and singing voice generation during the content-style modeling stage (Figure 3): (1) Implicit Prosody Learning (IPL): the AR model generates content-style tokens solely from text input, implicitly inferring prosodic information through in-context learning. This strategy is commonly adopted in zero-shot TTS systems [audiolm, seedtts, cosyvoice, vevo, maskgct]. (2) Explicit Prosody Learning (EPL): in addition to the text input, prosody tokens are explicitly provided, allowing the AR model to predict content-style tokens conditioned on both text and prosody information. Notably, during the joint training, we do not restrict IPL to speech data and EPL to singing voice data. Instead, we treat all samples equally: for each training sample, we randomly choose between IPL and EPL with equal probability. We assume that this design better unifies the training patterns for speech and singing voice.
Given a speech or singing voice sample u{u}, we denote the prosody and content-style tokens as $\bm{Q}_{p}({u})$ and $\bm{Q}_{cs}({u})$, respectively. The input sequences to the transformer for IPL and EPL are shown in Equation (3) and (4), where $\bm{I}_{ipl}$ and $\bm{I}_{epl}$ represent the text instruction prefixes for IPL and EPL, $\bm{T}$ denotes the text of u{u}, and $\langle|\cdot|\rangle$ indicates special tokens:
$\textbf{IPL}:[\bm{I}_{ipl},\bm{T},\langle|\text{start-of-cs}|\rangle,\bm{Q}_{cs}({u}),\langle|\text{end-of-cs}|\rangle],$ (3)
$\begin{split}\textbf{EPL}:[\bm{I}_{epl},\bm{T},\langle|\text{start-of-p}|\rangle,\bm{Q}_{p}({u}),\langle|\text{end-of-p}|\rangle,\\ \langle|\text{start-of-cs}|\rangle,\bm{Q}_{cs}({u}),\langle|\text{end-of-cs}|\rangle].\end{split}$ (4)
Specifically, the text instruction prefixes for IPL ($\bm{I}_{ipl}$) and EPL ($\bm{I}_{epl}$) are defined as: "User will provide you with a text. Please vocalize it with natural expression." and "User will provide you with a text. Please first generate a good prosodic instruction, then vocalize the text based on it." respectively. We perform next token prediction only on the final segment $\langle|\text{start-of-cs}|\rangle,\bm{Q}_{cs}({u}),\langle|\text{end-of-cs}|\rangle]$ for both IPL and EPL.
During the acoustic modeling stage, we employ an FM transformer to convert the content-style tokens into Mel spectrograms, conditioned on a timbre reference. This process follows the general paradigm used in speech generation studies [seedmusic, cosyvoice, vevo, glm4-iclr], with the distinction that our approach leverages the proposed unified content-style tokenizer and is jointly trained on both speech and singing voice data. After obtaining the Mel spectrograms, we use a Vocos-based vocoder [vocos], which is also trained on both speech and singing voice, to reconstruct the final waveform.
III-C Multi-Objective Alignment
After pre-training, we observe that the AR model exhibits good versatility across diverse tasks (Section V). Nevertheless, its stability remains suboptimal, particularly in its fundamental capabilities of text- and prosody-following, which can be further elicited and enhanced. Furthermore, in downstream applications, the model may encounter out-of-pretraining data such as instrumental sounds. For instance, in the context of SVS, the model needs to accept prosody tokens extracted from MIDI-rendered sounds for melody control (Figure 1).
Motivated by the above, this study further employs post-training to enhance the model’s controllability for text and prosody, as well as to improve its robustness to diverse audio data. We propose a multi-objective post-training task that jointly aligns both intelligibility and prosody similarity (Figure 4). We believe that this joint alignment helps prevent the model from over-optimizing a single objective (as evidenced in Section V-F).
Intelligibility Reward To align intelligibility, we adopt a recent zero-shot TTS intelligibility preference corpus, INTP [intp], which contains diverse intelligibility preference pairs constructed through comparisons between multiple models as well as manually designed intelligibility perturbations, to train an intelligibility reward model. Specifically, we employ the commonly used Bradley-Terry (BT) approach [bt-rm, instructgpt, RLHF-anthoropic] for reward model training. Given an intelligibility preference dataset $\mathcal{D}=\{(t,a^{w},a^{l})\}$, where $a^{w}$ and $a^{l}$ represent the positive and negative audio outputs with respect to the target text $t$, we parametrize an intelligibility reward model $r_{\phi}(t,a)$, treat the problem as binary classification, and optimize the negative log-likelihood loss:
@@ III-C Multi-Objective Alignment
where $r_{\phi}(t,a)$ is initialized from the pre-trained model. We follow the pre-training and use both EPL and IPL for training this reward model. During inference, we can input it a $\bm{T},\bm{Q}_{cs}]$ or $\bm{T},\bm{Q}_{p},\bm{Q}_{cs}]$ sequence to obtain an intelligibility reward.
Prosody Similarity Reward To achieve prosodic alignment, we establish the following protocol to evaluate the policy model’s prosody-following capability: (1) Suppose we have a singing voice dataset with MIDI annotations. Given a singing voice sample u{u}, we use its MIDI-rendered instrumental music to extract prosody tokens as control conditions, denoted as $\bm{Q}_{p}$; (2) Let the policy model perform a completion conditioned on $\bm{Q}_{p}$ and $\bm{T}$ (where $\bm{T}$ can be any text), and denote the generated content-style tokens as $\hat{\bm{Q}}_{cs}$; (3) Using the decoder of the content-style tokenizer (Section III-A), we reconstruct the chromagram features $\hat{\bm{x}}$ from $\hat{\bm{Q}}_{cs}$, and compute the cosine similarity between $\hat{\bm{x}}$ and the ground-truth chromagram features $\bm{x}$ of the singing voice u{u}. [^3] Note that in step (3), we could also use the flow-matching model and vocoder to perform inference and extract $\hat{\bm{x}}$ from the predicted waveform, but for training efficiency, we adopt the current approach.
Multi-Objective Optimization We apply the GRPO algorithm [grpo] to optimize the proposed multi-objective alignment task. Assuming the number of completions per prompt is $K$, for each prompt, after obtaining the intelligibility rewards ($\bm{r}_{int}$) and prosody similarity rewards ($\bm{r}_{pro}$) for all $K$ completions, we apply group normalization to both rewards separately. The sum of these normalized rewards serves as the final outcome supervision:
@@
III-D Inference-Time Controllability
@@
Text, Prosody, Style, and Timbre Control After obtaining the pre-trained (or post-trained) models, we can perform inference following the pipeline illustrated in Figure 1, by utilizing different combinations of text, prosodic source, style reference, and timbre reference to achieve versatile synthesis, conversion, and editing tasks for both speech and singing voice.
During the content-style modeling stage, for example, when the prosodic source is not used, the inference paradigm is identical to classic zero-shot TTS [valle, seedtts, vevo, cosyvoice]. As illustrated in the TTS task in Figure 1, the process can be represented as $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$], where the style signals in $\boldsymbol{\bm{\hat{cs}}}$ are prompted by $\boldsymbol{\bm{cs}}$ and generated through the in-context learning capabilities of the pre-trained model. Particularly, when both prosodic source and style reference are utilized as inputs, the inference paradigm takes the form $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{cs}}}$]. In this context, for $\boldsymbol{\bm{\hat{cs}}}$, we expect $\boldsymbol{\bm{p}}$ to provide coarse-grained prosody contour signals, while $\boldsymbol{\bm{cs}}$ contributes fine-grained style details (e.g., the vibrato patterns for singing). We will discuss this further through our experiments on style-converted singing voice conversion (Section V-C).
During the acoustic modeling stage, we can introduce timbre control by providing a timbre reference. Concretely, we concatenate the content-style tokens generated by the AR model with those extracted from the timbre reference, and jointly condition the FM model on the reference’s Mel spectrogram, i.e., $\boldsymbol{\bm{cs}}$, $\boldsymbol{\bm{mel}}$, $\boldsymbol{\bm{\hat{cs}}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{mel}}}$] of Figure 1. This dual conditioning guides the model to synthesize the target utterance with the linguistic, prosodic, and stylistic content from the AR output, while adopting the timbre characteristics embedded in the reference. It is noteworthy that we can also use only the FM model to conduct style-preserved VC or SVC tasks. Specifically, the AR-generated content-style tokens can be replaced with those extracted directly from a source audio, i.e., $\boldsymbol{\bm{cs}}$, $\boldsymbol{\bm{mel}}$, $\boldsymbol{\bm{cs}}$] $\rightarrow$ $\boldsymbol{\bm{\hat{mel}}}$] of Figure 1.
@@
Duration Control Some Non-AR based speech generation models have often featured the ability to control the total duration of generated speech [maskgct, f5tts], while AR-based models are generally considered less capable of controlling output duration [audiolm, seedtts, cosyvoice, vevo, indextts2]. Interestingly, however, we discover that Vevo2 can effectively control the duration of generated results through the use of prosody tokens as input. This capability stems from our tokenizers’ fixed frame rates (6.25 Hz and 12.5 Hz), which maintain a consistent 2:1 ratio between content-style and prosody token sequences during training. Consequently, during inference, we can control the duration of the generated output by applying simple linear scaling transformations to the chromagram features of the prosodic source, thereby modifying the length of prosody tokens. Such control mechanism is detailed in Section V-D.
@@
Pitch Region Control For voice conversion (VC) or singing voice conversion (SVC) tasks, due to potential significant differences in timbre types [vocalset, singingvoice-science] between source and reference speakers, we often need to adjust the pitch region (i.e., apply pitch shift) of the source waveform during inference to achieve higher speaker similarity [svcc-2023, amphion-svc, neucosvc]. In Vevo2, we discover that pitch region control can also be achieved: by shifting the pitch of the source waveform before prosody token extraction, we can effectively control the pitch region of the generated speech or singing voice. Our experiments demonstrate that this approach is effective both when using only the FM stage and when combining AR and FM stages. Such control mechanism is detailed in Section V-E.
IV Experimental Setup
IV-A Data and Metrics
Training Data For pre-training, we use Emilia [emilia] as the speech data (101K hours), and follow SingNet’s pipeline [singnet] to prepare 7K hours of singing voice data (source-separated from in-the-wild songs from the Internet). For post-training, we utilize INTP [intp] (which contains 250K preference speech pairs) to train the intelligibility reward model. For the proposed multi-objective alignment task, we employ three types of prosody prompts: (1) speech data: 20K positive samples randomly selected from INTP, (2) singing voice data: 20K samples from M4Singer [m4singer], and (3) instrumental sounds: rendered from MIDIs in M4Singer, covering 16 instruments across five categories: pianos, strings, woodwinds, brass, and folk instruments.
Evaluation Data We consider various settings to construct the evaluation set: (1) For regular speech, we utilize SeedTTS test sets [seedtts], which is widely adopted in the TTS field [maskgct, cosyvoice2, f5tts]. (2) For expressive speech, we employ Genshin-Voice [genshin_voice_dataset], which consists of professional voice acting recordings from the game Genshin Impact. (3) For singing voice, we use GTSinger [gtsinger] and SingStyle111 [singstyle111], which encompass a wide range of singing techniques, styles, and timbre types. (4) Additionally, for two unique tasks: humming-to-singing and instrument-to-singing, we utilize humming [humtrans] and instrumental sound including flute [flute-dataset] and violin [violin-dataset] for the evaluation. All evaluations are conducted using target text in both English and Chinese languages.
Evaluation Metrics For the objective metrics, we evaluate the intelligibility (WER, $\downarrow$), speaker similarity (SIM, $\uparrow$), F0 correlation (FPC, $\uparrow$[svcc-2023, amphion-svc, vevo]. For WER, we employ Whisper-large-v3 [whisper] for English texts, and Paraformer-zh [paraformer, funasr] for Chinese texts. For SIM, we compute the cosine similarity between the WavLM TDNN [wavlm] speaker embeddings of generated samples and the reference speeches. For subjective metrics, we employ Comparative Mean Opinion Score (CMOS, rated from -2 to 2, $\uparrow$) to evaluate naturalness (N-CMOS) and similarity in speaker, prosody, melody, and style (SS-CMOS, PS-CMOS, MS-CMOS, and Style-CMOS). Besides, we use Melody-MOS to evaluate the melody-following ability (rated from 1 to 3, corresponding to “unable to follow”, “roughly following the melody contour”, and “completely following all melody details”). Detailed backgrounds of subjects and definitions of all subjective metrics are provided in the supplementary materials.
IV-B Implementation Details
Prosody Tokenizer and Content-Style Tokenizer This paper processes and trains on speech and singing voice data with a sampling rate of 24 kHz. For chromagram feature extraction, we employ Librosa’s implementation [^4] https://librosa.org/doc/main/generated/librosa.feature.chroma_stft.html with the following parameters: number of FFT points set to 1920, hop length of 480, window size of 1920, and 24 chroma bins. For Whisper hidden features [whisper], we utilize the encoder output from the pre-trained whisper-medium model [^5] https://huggingface.co/openai/whisper-medium. We adopt the VQ-VAE implementations from Amphion [^6] https://github.com/open-mmlab/Amphion/tree/main/models/vc/vevo [amphion, amphion_v0.2], which serves for MaskGCT [maskgct] and Vevo [vevo]. Our only modification is the addition of a CNN-based downsampling layer before the encoder input and a CNN-based upsampling layer after the decoder output. The extracted chromagram features and Whisper features are both at 50 Hz. For the prosody tokenizer, the down(up)-sampling ratio is 8, resulting in a token rate of 6.25 Hz, with a codebook vocabulary size of 512. For the content-style tokenizer, the down(up)-sampling ratio is 4, yielding a token rate of 12.5 Hz, with a codebook vocabulary size of 16,384.
All models in this study use AdamW [adam, adamw] as the optimizer. For the two tokenizers, we use a peak learning rate of 1e-4, with linear warm-up for 10K steps followed by decay over the remaining training period. Both tokenizers are trained for 300K updates. The prosody tokenizer and content-style tokenizer contain 38M and 44M parameters, respectively.
Content-Style Modeling During the content-style modeling stage, we initialize the AR model using Qwen2.5-0.5B [^7] https://huggingface.co/Qwen/Qwen2.5-0.5B [qwen2.5] and expand its original vocabulary to accommodate our proposed prosody tokens and content-style tokens. The total parameter count of the AR model is 509M. For the BPE-based text tokenizer, we adopted the same tokenization strategy as CosyVoice 2, which masks out one-to-many tokens to prevent excessive pronunciation length and mitigate corner cases caused by data sparsity [cosyvoice2]. During pre-training, we use a peak learning rate of 5e-4, with linear warm-up for 32K steps followed by decay over the remaining training period. The model is pre-trained for 500K updates.
Acoustic Modeling During the acoustic modeling stage, we adopted the flow-matching (FM) transformer architecture from Vevo [vevo]. The transformer consists of 16 layers with 16 attention heads and a hidden dimension of 1024. The FM model contains 363M parameters in total. To enhance training efficiency, we incorporated the REPA strategy [repa] and enforced alignment between the hidden output of the transformer’s 5th layer and the W2v-BERT 2.0 features [^8] https://huggingface.co/facebook/w2v-bert-2.0 [w2v-bert, maskgct] during the training process. We use a peak learning rate of 7.5e-5, with linear warm-up for 32K steps followed by decay over the remaining training period. The FM model is trained for 700K updates. To develop a unified vocoder for both speech and singing voice, we fine-tuned Vevo’s speech vocoder [^9] https://huggingface.co/amphion/Vevo [vevo] using our training data (comprising 101K hours of speech and 7K hours of singing voice data). The vocoder, which is based on the Vocos architecture [vocos], contains 255M parameters and underwent fine-tuning for 572K updates.
Post-training During the post-training, for each training iteration with the 20K M4Singer samples, we randomly selected one instrument and rendered the corresponding audio using pretty_midi [^10] https://github.com/craffel/pretty-midi. We use the original text of the sample and extract prosody tokens from these rendered instrumental sounds. We adopt the training framework of TRL [^11] https://huggingface.co/docs/trl/main/grpo_trainer. We use the learning rate of 5e-6, the batch size of 3, the rollout number of 8 and the KL coefficent is 0.1. The model is post-trained for 9K updates.
V Results and Analysis